OpenStudy (anonymous):

describe the vector \(\large \frac{1}{3}\mathbb{u}+ \frac{1}{3}\mathbb{v}+ \frac{1}{3}\mathbb{w}\)

OpenStudy (perl):



OpenStudy (anonymous):

\(\large \mathbb{u},\mathbb{v},\mathbb{w}\) are independent vectors in 3D

OpenStudy (perl):